In this balanced chemical equation, 2 C18H30O2 + 49O2 à 36 CO2 + 30 H2O, answer the following questions below:

a. How many atoms of O are found on both sides of the equation? --

b. How many atoms of H are found on both sides of the equation? --

c. What is the mole ratio between O2 and CO2? --

d. What is the mole ratio between C18H30O2 and H2O? --

e. What type of reaction is this?
Answer #1

balanced chemical equation :  2 C18H30O2 + 49O2  \rightarrow 36 CO2 + 30 H2O

a. 'O ' atoms : 102 atoms on both sides of equation :

(2*2 + 49*2 = 102 ) and ( 36*2 + 30 *1 = 102 )

b. 'H' atoms :60 atoms both sides of equation :

( 2*30 = 60) and ( 30*2 = 60)

c. Mole ratio between O2 and CO2 :

O2   moles = 49 ;  CO2 moles = 36

O2 and CO2 mole ratio = 49 / 36 = 1.36

d. Mole ratio between  C18H30O2 and H2O :

C18H30O2 moles = 2 H2O moles = 30

C18H30O2 and H2O mole ratio = 2/30 = 0.067

e. This oxidation reaction.